A town has a population of 14000 and grows at 2.5% every year. What will be the population after 5 years, to the nearest whole number?


1
Expert's answer
2021-02-01T07:11:58-0500

initial population = 14000

growth rate = 2.5% every year

so after one year population is = 14000+14000(2.5/100) = 14350

after 2 year population is = 14350+14350(2.5/100) = 14709

this will continue till 5 years so reduce it in a formula as

"pop(n)= pop(initial) *(1+\frac{rate per year}{100})^n"


where,


pop(n) = population after n years.


pop(initial) = initial population.


n = number of years


so, population after n years with growth rate 2.5% is


"pop(5) = 14000(1+\frac{2.5}{100})^5"


"pop(5) = 14000(1.025)^5"

pop(5) = 15839.74


thus, population after 5 years is 15840.